# DHGS: Decoupled Hybrid Gaussian Splatting for Driving Scene

### v2

@@ Introduction
where $i$ and $j$ satisfying $\tilde{M}^{i,j}\not=0$, $\tilde{M}$ refers to the banded boundary generated by binary mask $M$, and the orange band at the bottom right corner of Figure 2 intuitively shows the distribution of elements in $\tilde{M}$. We follow Plenoxels (Fridovich-Keil et al. 2022) and use a total variation loss $\mathcal{L}_{\text{tv}}$. The collective impact of the aforementioned loss functions enables DHGS to generate high-quality renderings that preserve a continuous geometric structure.
@@
Datasets and Metrics. We experiment on the publicly available urban scene datasets Waymo (Sun et al. 2020), 32 static scenes are adopted as the same as EmerNeRF (Yang et al. 2023a). We employ multiple metrics to evaluate the performance of the proposed method and baselines. PSNR, SSIM, and LPIPS metrics are utilized to validate the quality of rendered images with ground truth. For novel views without ground truth, we introduce FID (Heusel et al. 2017) to evaluate the image quality. The averaged metrics across all clips are calculated as comparative values.
Baselines. We compare our method with state-of-the-art and publicly available works. 3DGS (Kerbl et al. 2023), 2DGS (Huang et al. 2024a), Gaussian Pro (Cheng et al. 2024), and Scaffold GS (Lu et al. 2024) are chosen as the comparison methods having been regarded for their outstanding performances.
@@
Implementation details. The initial point cloud is randomly downsampled to 600,000 points. All methods are trained for 60,000 iterations. For the pre-trained SDF network, we employ an 8-layer MLP and train it for 20,000 iterations. The normals of surface points are computed using SVD with 50 neighbors, and the weights for the normal loss and the eikonal loss of the SDF network are set to 0.1 and 0.01. The weights $\lambda_{\text{tran}}$, $\lambda_{\text{sdf}}$, $\lambda_{\text{cons}}$, $\lambda_{\text{tv}}$, $\lambda_{d}$ and $\lambda_{n}$ of the overall training loss are set to 0.1, 1.0, 0.04, 0.1, 0.1 and 0.1. We set the densification intervals for the Gaussian ellipsoids in the environment model and road model to 200 and 300 respectively. The learning rate for the anchor position in the environment model is set as 0.00016, the remaining learning rates are kept consistent with the original settings. For each clip, we utilize all the front, left-front, and right-front views of the entire sequence for training, with a 4:1 split ratio between the training and testing sets. The training set is referred to as the reconstruction views, while the testing set is called the novel views. We further apply unrestricted transformations on camera poses to validate the potential of the model on data-driven simulation, which is referred to as free-view novel view synthesis.
@@
We benchmark our method against established baselines in scene reconstruction, novel view synthesis, and free-view novel view synthesis that stands for notable viewpoints transfer. All methods employ a consistent set of reconstruction views for training, while the remainder are employed as novel views for testing in each clip. The qualitative and quantitative results will be examined in the subsequent sections, with an ablation study included for analysis.
@@
Scene Reconstruction on the Waymo Dataset. For reconstruction views, we calculate the mean PSNR, SSIM, and LPIPS of the rendered images across 32 clips, as shown in Table 1. Our method outperforms all the mainstream methods: including 3DGS, 2DGS, Gaussian Pro, and Scaffold GS. The left part of Figure 4 displays the rendered images of the reconstruction views from our method and its competitors. It can be noticed that our method achieves better reconstruction quality with finer details than competitive methods, from which the road area rendered by our method contains richer texture information.
@@
Novel View Synthesis on the Waymo dataset. We assess the aforementioned methods on the novel views. Table 1 illustrates the validation results, demonstrating that our method surpasses most rival methods, such as 3DGS, 2DGS, and Gaussian Pro, and achieves comparable results to Scaffold GS. Notably, Scaffold GS achieves the best performance in terms of PSNR and SSIM. As depicted in the right part of Figure 4, blurriness appears in some non-road areas (highlighted by red boxes) for 3DGS, 2DGS, and Gaussian Pro. In contrast, our method restores more road details (lane lines within the red boxes) thanks to the imposition of geometric constraints on the road, whereas other methods such as 2DGS, Gaussian Pro, and Scaffold GS fail to preserve these subtleties. Despite novel views not being part of the training set, their camera poses variations are minimal compared to trained views, which can be simply concluded as the interpolation of trained views. Such minor variations in novel views do not sufficiently reflect the geometric representation capability of the model.
@@
Free-view Novel View Synthesis on the Waymo dataset. To further underscore the geometric consistency of our approach, we evaluate the trained model on free-view novel views with more pronounced camera pose variations. We adjust the poses of the cameras relative to the vehicle, mimicking various observation perspectives for different vehicle types with diverse camera positions. In this paper, we outline four types of transformations to the camera poses, with detailed configurations provided in the supplementary material. Given the absence of ground truth for these novel views, we adopt the FID metric, adept at gauging the feature discrepancy between the original input views and the free-view novel views. The comparisons of FID for free-view novel views are also presented in Table 1, where our method achieves the highest performance compared to other methods in all four configurations with a significant margin of superiority. The qualitative rendering results, depicted in Figure 5, reveal that free-view rendered images produced by other methods suffer from extensive artifacts and lane line breakage. In contrast, our method consistently captures environmental and road details, including lane lines, even amidst substantial camera pose shifts. The results further validate that the regularization constraints and the depth-ordered hybrid rendering strategy we proposed facilitate the acquisition of an optimal geometric structure, thereby showcasing the robustness of our approach in free-view novel view synthesis. More visual comparisons can be found in supplemental materials.
@@
Ablation Study.
To assess the efficacy of our proposed regularization components, we perform an ablation study, systematically omitting the SDF, transmittance, and consistency losses, denoted as ”w/o SDF”, ”w/o tran”, and ”w/o cons”. The quantitative results for scene reconstruction, novel view synthesis, and free-view novel view synthesis are consolidated in Table 2. For scene reconstruction and novel view synthesis, the absence of the SDF loss leads to a decline in performance metrics, whereas the exclusion of transmittance loss or consistency loss has a minimal impact. Conversely, for free-view novel view synthesis, the removal of any loss escalates the FID score, signifying a greater divergence between the novel views and the originals. To more intuitively illustrate the specific contributions of these components, we present the visualization results of free-view novel views synthesis in Figure 6. Specifically, the SDF loss ensures the fidelity of road geometry, the consistency loss guarantees a seamless connection between the environment model and the road model. The transmittance loss ensures precise alignment of the rendering areas of both models with the 2D segmentation, thus averting rendering artifacts. Therefore, each regularization term is integral to our DHGS framework.
@@ Conclusions
In this paper, we propose a novel reconstruction method called DHGS, which aims to elevate the rendering quality of novel view synthesis for static driving scenes. DHGS leverages two decoupled models to represent the road and environment separately, which are then merged via a pixel-level hybrid renderer, facilitated by our innovative depth-ordered hybrid rendering strategy. Additionally, consistency loss and transmittance loss are proposed to constrain the basal models, allowing for the generation of continuous and consistent rendered images. Moreover, we develop a road surface regularization term grounded in SDF to ensure better geometric consistency of the road surface, guaranteeing stable image quality as perspective changes drastically. Through extensive testing on the Waymo dataset, our method reveals the state-of-the-art performance in both reconstruction and novel view synthesis of our method.
Acknowledgments
This work would thank Bo Zheng and Fan Jin for providing the necessary help in the preparation of training data.
@@ A. Implementation Details
To deal with the floaters phenomenon in new perspectives, we devise a straightforward yet effective strategy. We construct a hemispherical virtual sky point cloud layer that envelops the LiDAR point clouds. The radius and center of this sky point cloud are determined by the scale and the average distribution of the LiDAR points, ensuring an accurate representation of the sky’s spatial characteristics. Additionally, we apply appropriate coloring to the sky point cloud to maintain its visual integrity, as depicted in Figure A1. Table A1 presents a comparative analysis between incorporating the sky sphere point cloud into 3DGS and not incorporating it. It is evident that the introduction of sky sphere pcd is able to improve the rendering quality, no matter in reconstruction or novel view synthesis. For fairness, we integrate this sky sphere point cloud mechanism into all comparative methods, such that the potential biases stemming from differences in sky handling can be eliminated.
@@
PCD wo sky sphere 25.63 0.8062 0.3265 67.52
PCD wi sky sphere 25.92 0.8089 0.3218 64.77
Table A1: The quantitative results of 3DGS with and without incorporating the sky sphere point cloud into the initial point cloud. The FID here is averaged on FIDS of all four sets.
@@
In addition to optimizing the road model, we also utilize a distance constraint mechanism for the environmental model. Rather than directly applying distance regularization to the offsets, we adopt a more restrictive approach, overlaying a sigmoid function on the offsets. This achieves effective boundary control over the offsets. Specifically, we implement the transformation using the following formula:
@@
where $x_{v}$ refers to the coordinate of the anchor point, $\mu_{k}$ refers to the center coordinates of the $k$-th neural Gaussian corresponding to the anchor. $l_{v}$ and $\mathcal{O}_{v}$ represent scaling factor and offset vector respectively. This design ensures that the offsets are strictly confined within a reasonable and finite range, thereby preventing the generation of neural Gaussian ellipsoids that are excessively distant from their anchor points. Consequently, the power of the environmental model to represent fine details is significantly enhanced.
@@
The total variation regularization $\mathcal{L}_{\text{tv}}$ (Rudin and Osher 1994) is applied in this study to refine the rendered pixel depths. The depth map can be generated through a rendering process similar to the one we calculate color:$D=\lambda_{r}D_{r}+\lambda_{e}D_{e}$, to maintain the continuity of $D$ near boundaries of two depth maps, the TV regularizer can be represented as:
@@
The SDF regularizer $\mathcal{L}_{\text{sdf}}$, inspired by NeuSG originally optimizing the positions and orientations of 3D Gaussians, is applied in this paper similarly to 2D Gaussians. Specifically, we leverage the gradient direction of the pre-trained SDF to guide the orientation updates of the 2D Gaussians, while simultaneously constraining their disk centers to align with the SDF. This approach guarantees a seamless integration which is intuitively illustrated in Figure A2.
@@
Table A2: The table illustrates four different parameter settings for free-view novel view synthesis, where Translation refers to the translation component, Rotation refers to the rotation component, the corresponding three numbers are the lengths of translation in meter, and the magnitudes of three Euler angles for rotation in degree.
@@ B. Experimental results
We apply four transformations to the pose of the camera relative to the vehicle, including rotations and translations. The motive is that transferring the visual data from the collecting vehicle to a new vehicle with a diverse sensor setup is frequently demanded in AD. We set four settings based on the camera positions of the Waymo dataset to simulate the camera displacements caused by changes in vehicle type. The detailed variation amplitudes for the front, front-left, and front-right camera under each transformation are outlined in Table A2, where Translation denotes the translation distances of the camera along the $x$, $y$, and $z$ axes in meter, and Rotation represents the rotation angles arorund the $yaw$, $pitch$, and $roll$ axes in degree. Figure A3 depicts the differences in the four transformed camera poses compared to the original Waymo configuration.
@@
$\text{E}_{\text{Scaffold GS}}$ + $\text{R}_{\text{2DGS}}$ 28.09 26.77 48.05 56.94 75.72 55.52
@@
Figure A4: Qualitative comparisons of different compound modes on the Waymo dataset, with top ($\text{E}_{\text{3DGS}}$ + $\text{R}_{\text{3DGS}}$), middle ($\text{E}_{\text{Scaffold GS}}$ + $\text{R}_{\text{3DGS}}$), and bottom ($\text{E}_{\text{Scaffold GS}}$ + $\text{R}_{\text{2DGS}}$). The highlighted regions demonstrate that our method achieves the best rendering quality.
@@
The qualitative result on the Waymo dataset.
@@
Comparative analysis of the compound mode of Gaussians.
We compare several distinct assembly ways of road and non-road models, as detailed in Table A3. Figure A4 illustrates qualitative comparisons of three compound modes. It can be concluded that leveraging Scaffold GS and 2DGS to represent the environment and road model severally harvests optimal performance. Benefiting from the robustness of Scaffold GS in changing views and the inherent superiority of flat 2DGS representing the road surface, DHGS achieves superior rendering results both in the road and environment areas, exhibiting fewer floaters on the road and capturing finer environmental details..

### v3

@@ Introduction
where $i$ and $j$ satisfying $\tilde{M}^{i,j}\not=0$, $\tilde{M}$ refers to the banded boundary generated by binary mask $M$. The orange band intuitively indicates the non-zero entries in $\tilde{M}$, as shown in the bottom right of Figure 2. We follow Plenoxels (Fridovich-Keil et al. 2022) and use a total variation loss $\mathcal{L}_{\text{tv}}$. The cumulative effect of the aforementioned loss functions allows DHGS to produce high-fidelity renderings that maintain a continuous geometric structure.
@@
Datasets and metrics. We experiment on the publicly available urban scene dataset Waymo (Sun et al. 2020), 32 static scenes are adopted as the same as EmerNeRF (Yang et al. 2023a). We employ multiple metrics to evaluate the performance of the proposed method and baselines. PSNR, SSIM, and LPIPS are utilized to validate the image quality for rendered images with ground truth. For novel views lacking ground truth, we apply FID score (Heusel et al. 2017) to assess the image quality. The average metrics across all clips are calculated as comparative values.
Baselines. We compare our method with the state-of-the-art and publicly available methods. 3DGS (Kerbl et al. 2023), 2DGS (Huang et al. 2024a), Gaussian Pro (Cheng et al. 2024), and Scaffold GS (Lu et al. 2024) are chosen for comparison, recognized for their exceptional performances.
@@
Implementation details. The initial point cloud is randomly downsampled to 600,000 points. Our method and the other comparison methods are trained for 60,000 iterations. For the pre-trained SDF network, we use an 8-layer MLP, training it for 20,000 iterations. The normals of surface points are derived using SVD with 50 neighborhoods. The weights for the normal and eikonal losses in the SDF network are carefully calibrated, set to 0.1 and 0.01, respectively. The weights $\lambda_{\text{tran}}$, $\lambda_{\text{sdf}}$, $\lambda_{\text{cons}}$, $\lambda_{\text{tv}}$, $\lambda_{d}$ and $\lambda_{n}$ of the overall training loss are set to $0.1$, $1.0$, $0.04$, $0.1$, $0.1$ and $0.1$. We set the densification intervals for the Gaussian primitives in the environment model and road model to $200$ and $300$ respectively. The learning rate for the anchor position in the environment model is set as $0.00016$, the remaining learning rates are kept consistent with the original settings. For each clip, we utilize all the front, left-front, and right-front views of the entire sequence for training, with a 4:1 split ratio between the training and testing sets. For simplicity, the images in the training set are referred to as the reconstruction views, while those in the testing set are called the novel views. We further apply unrestricted transformations to camera poses to test the robustness of the model in data-driven simulation, termed as free-view novel view synthesis.
@@
We benchmark our method against established baselines in scene reconstruction, novel view synthesis, and free-view novel view synthesis which signifies notable viewpoint transfer. All methods employ a consistent set of reconstruction views for training, while the remainder are employed as novel views for testing in each clip. Qualitative and quantitative results are shown in the following sections, including an ablation study for detailed analysis.
@@
Scene Reconstruction on the Waymo dataset. For reconstruction views, we calculate the mean PSNR, SSIM, and LPIPS of the rendered images across 32 clips, as shown in Table 1. Our method outperforms all the comparative methods: including 3DGS, 2DGS, Gaussian Pro, and Scaffold GS. The left part of Figure 4 displays the rendered images of the scene reconstruction from our method and its competitors. It can be noticed that our method delivers superior reconstruction quality with finer details, particularly in the road area, where our rendering exhibits richer texture information.
@@
Novel View Synthesis on the Waymo dataset. We assess the aforementioned methods on the novel view synthesis. Table 1 illustrates the validation results, demonstrating that our method surpasses most rival methods, such as 3DGS, 2DGS, and Gaussian Pro, and achieves comparable results to Scaffold GS. Notably, Scaffold GS leads in PSNR and SSIM metrics. As depicted in the right part of Figure 4, blurriness that is highlighted by red box appears in some non-road areas for 3DGS, 2DGS, and Gaussian Pro. In contrast, our method restores more road details which can be seen in lane lines within the red boxes, owing to the imposition of geometric constraints on the road. Whereas, other methods such as 2DGS, Gaussian Pro, and Scaffold GS fail to preserve these subtleties. Despite novel views not being part of the training set, their camera pose variations are minimal compared to trained views, suggesting they can be considered as interpolations of the trained views. Such minor variations in novel views do not sufficiently reflect the capability of geometric representation of the model.
@@
Free-view Novel View Synthesis on the Waymo dataset. To further underscore the geometric consistency of our approach, we evaluate the trained model on free-view novel views with more pronounced camera pose variations. We adjust the poses of the cameras relative to the vehicle, mimicking various observation perspectives for different vehicle types with diverse camera positions. This paper details four types of camera pose transformations, with specific configurations available in the supplementary material. Given the absence of ground truth for these novel views, we adopt the FID score, which excels at measuring the feature discrepancy between the original and the free-view novel views. The comparisons of FID for free-view novel views are also presented in Table 1, where our method achieves the highest performance compared to other methods in all four configurations with a significant margin of superiority. The qualitative rendering results, depicted in Figure 5, reveal that free-view rendered images produced by other methods suffer from extensive artifacts and lane line breakage. In contrast, our method consistently captures environment and road details, including lane lines, even with substantial variations in camera poses. The results further confirm that the regularization constraints and the depth-ordered hybrid rendering strategy we proposed facilitate the acquisition of an optimal geometric structure, thereby showcasing the robustness of our approach in free-view novel view synthesis. More visual comparisons can be found in supplemental materials.
@@
Ablation Study
To assess the efficacy of our proposed regularization components, we perform an ablation study, systematically omitting the SDF, transmittance and consistency losses, denoted as ”w/o SDF”, ”w/o tran”, and ”w/o cons”. The quantitative results for scene reconstruction, novel view synthesis, and free-view novel view synthesis are consolidated in Table 2. For scene reconstruction and novel view synthesis, the absence of the SDF loss leads to a decline in performance metrics, whereas the exclusion of transmittance loss or consistency loss has a minimal impact. Conversely, for free-view novel view synthesis, the removal of any loss escalates the FID score, signifying a greater divergence between the novel view and the originals. To more intuitively illustrate the specific contributions of these components, we present the visualization results of free-view novel view synthesis in Figure 6. Specifically, the SDF loss ensures the fidelity of road geometry, while the consistency loss guarantees a smooth integration between the environment and road models. The transmittance loss is crucial for the precise alignment of the rendering regions of both models with the 2D segmentation, thereby averting rendering artifacts. Therefore, each regularization term is integral to our DHGS framework.
@@ Conclusions
In this paper, we propose a novel reconstruction method called DHGS, which aims to elevate the rendering quality of novel view synthesis for static driving scenes. DHGS leverages two decoupled models to represent the road and environment separately, which are then merged via a pixel-level hybrid renderer, facilitated by our innovative depth-ordered hybrid rendering strategy. Additionally, consistency loss and transmittance loss are proposed to constrain the basal models, enabling the generation of continuously consistent rendered images. Moreover, we develop a road surface regularization term grounded in SDF to ensure better geometric consistency of the road surface, guaranteeing stable image quality as perspective changes drastically. Through extensive testing on the Waymo dataset, our method demonstrates the state-of-the-art performance in reconstruction and free-view novel view synthesis.
@@ A. Implementation Details
To address the floaters phenomenon in new perspectives, we devise a straightforward yet effective strategy. We construct a hemispherical virtual sky point cloud layer that envelops the LiDAR point cloud. The radius and center of this sky point cloud are determined by the scale and the average coordinates of the LiDAR points, ensuring an accurate representation of the spatial properties within the sky. Additionally, we apply appropriate coloring to the sky point cloud to maintain its visual integrity, as depicted in Figure A1. Table A1 presents a comparative analysis between incorporating the sky sphere point cloud into 3DGS and not incorporating it. It is evident that the introduction of sky sphere pcd is able to improve the rendering quality, no matter in reconstruction or novel view synthesis. For fairness, we integrate this sky sphere point cloud mechanism into all comparative methods, such that the potential biases stemming from differences in sky handling can be eliminated.
@@
PCD w/o sky sphere 25.63 0.8062 0.3265 67.52
PCD w/ sky sphere 25.92 0.8089 0.3218 64.77
Table A1: The quantitative results of 3DGS with and without incorporating the sky sphere point cloud into the initial point cloud. The FID here is the average based on the FIDs of all four sets.
@@
In addition to optimizing the road model, we also utilize a distance constraint mechanism for the environment model. Rather than directly applying distance regularization to the offsets, we adopt a more restrictive approach, overlaying a sigmoid function on the offsets. This approach ensures more stringent control over the boundaries. Specifically, we implement the transformation using the following formula:
@@
where $x_{v}$ refers to the coordinate of the anchor point, $\mu_{k}$ refers to the center coordinates of the $k$-th neural Gaussian corresponding to the anchor. $l_{v}$ and $\mathcal{O}_{v}$ represent the scaling factor and offset vector respectively. This design ensures that the offsets are strictly confined within a reasonable and finite range, thereby preventing the generation of neural Gaussian ellipsoids that are excessively distant from their anchor points. Consequently, the power of the environment model to represent fine details is significantly enhanced.
@@
The total variation regularization $\mathcal{L}_{\text{tv}}$ (Rudin and Osher 1994) is applied in this paper to refine the rendered pixel depths. The depth map can be generated through a rendering process similar to the one we calculate the depth map: $D=\lambda_{r}D_{r}+\lambda_{e}D_{e}$, to maintain the continuity of $D$ near the boundaries of two depth maps, the TV regularizer can be represented as:
@@
The SDF regularizer $\mathcal{L}_{\text{sdf}}$, inspired by NeuSG (Chen, Li, and Lee 2023) originally optimizing the positions and orientations of 3D Gaussians, is applied in this paper similarly to 2D Gaussians. Specifically, we leverage the gradient direction of the pre-trained SDF to guide the orientation updates of the 2D Gaussians, while simultaneously constraining their disk centers to align with the SDF. This approach guarantees a seamless integration, which is intuitively illustrated in Figure A2.
@@
Table A2: The table illustrates four different parameter settings for the free-view novel view synthesis, where Translation refers to the translation component, Rotation refers to the rotation component, the corresponding three numbers are the lengths of translation in meters, and the magnitudes of three Euler angles for rotation in degree.
@@ B. Experimental results
We apply four transformations to the pose of the camera relative to the vehicle, including rotations and translations. The motive is that transferring the visual data from the collecting vehicle to a new vehicle with a diverse sensor setup is frequently demanded in AD. We set four settings based on the camera positions of the Waymo dataset to simulate the camera displacements caused by the changes in vehicle type. The detailed variation amplitudes for the front, front-left, and front-right camera under each transformation are outlined in Table A2, where Translation denotes the translation distances of the camera along the $x$, $y$, and $z$ axes in meters, and Rotation represents the rotation angles around the $yaw$, $pitch$, and $roll$ axes in degree. Figure A3 depicts the differences in the four transformed camera poses compared to the original Waymo configuration.
@@
Ours 28.09 26.77 48.05 56.94 75.72 55.52
@@
Figure A4: Qualitative comparisons of different compound modes on the Waymo dataset, with top ($\text{E}_{\text{3DGS}}$ + $\text{R}_{\text{3DGS}}$), middle ($\text{E}_{\text{Scaffold GS}}$ + $\text{R}_{\text{3DGS}}$), and bottom (Ours). The highlighted regions demonstrate that our method achieves the best rendering quality.
@@
The qualitative results on the Waymo dataset.
@@
Comparative analysis of the compound modes of Gaussians.
We compare several distinct assembly ways of road and non-road models, as detailed in Table A3 subsequently. Figure A4 illustrates qualitative comparisons of three compound modes. It can be concluded that leveraging Scaffold GS and 2DGS to represent the environment and road model severally harvests the optimal performance. Benefiting from the robustness of Scaffold GS in changing views and the inherent superiority of flat 2DGS representing the road surface, DHGS achieves superior rendering results both in the road and environment areas, exhibiting fewer floaters on the road and capturing finer environment details.